# Semantic-Aware Resource Management for C-V2X Platooning via Multi-Agent Reinforcement Learning

### v1

@@ III-A Modeling of the System
State space: The state of the $n$th PL at time slot $t$ consists of the following components: the instantaneous channel information of the $n$th PL and the BS $h_{n,BS}^{t}[k]$, the instantaneous channel information of the $n$th PL and his followers $h_{n,m}^{t}[k]$, the previous interference from other platoons $I_{T,BS}^{t}[k]$ and $I_{I,BS}^{t}[k]$, the residuals transmitted by the intra-platoon communication loads $B_{s,n}^{t}$, and thus, the state space can be expressed as:
$s_{n}^{t}=\left[h_{n,BS}^{t}[k],h_{n,m}^{t}[k],I_{T,BS}^{t}[k],I_{I,BS}^{t}[k],B_{s,n}^{t}\right].$ (23)
Action space: As mentioned before, PL contains several actions, $\beta_{n,k}$ is used to select the sub-channel for communication, inter or intra communication $\rho_{n,k}$, the power level of the communication $p_{T,n}[k]$ and $p_{I,n}[k]$ and the sentence lengths corresponding to V2I communication, V2V communication ($u_{m}^{T}$ and $u_{m}^{I}$). Note that when the PL selects inter communication, it has only one sentence length to select; when the PL selects intra communication, there will be $M$ followers of sentence lengths to select. Thus, the action space of agent $k$ at time slot $t$ is defined as:
@@
Reward function: in our proposed scheme, there are two aspects of rewards for each agent, one is the global reward reflecting the cooperation among agents, and the other is the local reward that is to help each agent to explore the optimal action. In DRL, rewards can be set flexibly, and a good reward can improve the performance of the system. Here, our goal is to optimize the agent $p$’s QoE level, as well as the corresponding payload transmission probability. Thus the local and global rewards of the $p$th PL at time slot $t$ are:
$\begin{array}[]{l}{r_{n,l}}=-{w_{1}}\frac{1}{{1+\exp(-\alpha(\sum\limits_{t=1}^{T}{\sum\limits_{m=1}^{M}{{\beta_{n,k}}}}\phi_{k}^{d}[m,t]-\frac{{{B_{s}}}}{{\Delta T}}))}}\\ {\rm{}}\qquad\qquad\qquad\qquad\qquad\qquad\quad+{w_{2}}Qo{E^{n}},\end{array}$ (25)
@@ III-B SAMRAMARL Algorithm
Algorithm 1 Modified MADDPG Algorithm
We propose a SAMRAMARL algorithm, which is designed to optimize both global system performance and individual agent behavior in a dynamic semantic communication environment. This method is particularly suited for vehicular platooning, where optimizing the transmission of semantically rich data is critical for enhancing the overall system performance and maintaining local adaptability. This algorithm two primary critics: a shared global critic that assesses states and actions across all agents to enhance global rewards, and individual local critics for each agent that evaluate rewards based on their local states and actions. The SAMRAMARL algorithm is designed to optimize both global system objectives and individual agent behavior in a dynamic, semantic communication environment, particularly tailored for vehicular platooning scenarios. It achieves this through a dual-critic framework, involving both global and local critic networks. The algorithm employs a multi-agent actor-critic approach, where each agent is equipped with its own actor and local critic networks for decision-making, while a pair of global critics guides the overall system optimization [57].
Specifically, we consider a vehicular environment with $P$ platoons and the policies for all agents are $\mathbf{\pi}=\{{\pi_{1}},{\pi_{2}},\cdots,{\pi_{N}}\}$. The agent $n$’s strategy $\pi_{n}$, Q-functions ${\mathop{\rm Q}\nolimits}_{{\phi_{n}}}^{k}$ and twin global critic Q-functions (${\mathop{\rm Q}\nolimits}_{\psi 1}^{{g_{1}}}$, ${\mathop{\rm Q}\nolimits}_{\psi 2}^{{g_{2}}}$) are parameterized by ${\theta_{n}}$, $\phi_{n}$, ${\psi_{1}}$ and $\psi_{2}$, respectively. The training process begins with an initial the environment simulator, configured with $P$ vehicular platoons to reflect a realistic multi-agent communication scenario. Twin global critic networks, $Q_{{\psi_{1}}}^{{g_{1}}}$ and $Q_{{\psi_{2}}}^{{g_{2}}}$ are initialized to evaluate joint state-action pairs across all agents. These global critics are responsible for assessing the performance of the entire system based on global rewards which is QOE. Alongside, each agent $n$ initializes its policy network ${\pi_{n}}$ for action selection and a local critic network $Q_{{\phi_{n}}}^{n}$ to evaluate the quality of its actions based on local observations.
@@
The detailed SAMRAMARL algorithm is described in the Algorithm 1. Since the modified MADDPG method is implemented as the actor-critic framework, denote the computational complexities for calculating gradients and updating parameters in the actor and critic networks being denoted as $G_{A}$, $G_{C}$, $U_{A}$, and $U_{C}$ respectively. Since the architecture of the target actor network and the target critic network is the same as the network structure of the two mentioned above, their complexities remain the same. Thus, the computational complexity of the modified MADDPG is calculated as $\mathcal{O}\left({{G_{A}}+{G_{C}}+2{U_{A}}+2{U_{C}}}\right)+{\text{ }}\mathcal{O}\left({2{G_{C}}+4{U_{C}}}\right)$.
@@ IV-B Performance Evaluation
To assess the performance of the proposed distributed semantic-aware multi-modal resource allocation framework (SAMRAMARL), we utilize QoE and SRS in V2X links as performance metrics. We compare SAMRAMARL against the following baseline methods:
@@
In traditional communication systems, semantic information is transmitted using bits rather than symbols explicitly designed for semantic representation. Although each bit can be loosely considered a semantic symbol, it generally conveys less semantic information compared to symbols employed in semantic transmission methods like DeepSC. The $QOE^{\prime}$ is defined similar to equation (15):
@@
where ${u_{m}^{T/I}}$ is defined as the transform factor, similar to equation (34), with unit of bits/word. For example, if a word consists of 5 letters and ASCII code is used for encoding, then ${u_{m}^{T/I}}$ is 40 bits/word. We assume that ${{\tilde{H}}_{SM/MM}}$ is fixed, indicating no bit errors in conventional communications.
@@
In Fig. 2, we analyze the reward evolution of various algorithms, including our SAMRAMARL framework. SAMRAMARL exhibits faster convergence and greater stability than $DDPG$ and $TD3$, attributable to its multi-agent structure. This decentralized decision-making allows each vehicle to respond independently to local conditions, optimizing resource allocation in dynamic environments. While $DDPG$ and $TD3$ also consider semantic awareness, their single-agent models struggle to adapt quickly to environmental changes. Specifically, $DDPG$ faces instability from a poor exploration-exploitation balance, resulting in fluctuating rewards, while $TD3$ converges more slowly than SAMRAMARL, highlighting the benefits of multi-agent systems. The distributed nature of SAMRAMARL fosters better collaboration, leading to quicker convergence and stability, particularly under rapidly changing semantic information. This approach effectively meets the varying demands of vehicles within a platoon, ensuring efficient spectrum sharing and resource allocation
Figure 3: Intra platoon Gap vs QOE
@@
Figure 4: Intra platoon Gap vs SRS
Fig. 4 presents the relationship between intra-platoon gap and SRS. The $DDPG_{N}O_{S}C$ curve shows a clear decline with increasing gaps, reflecting increased signal attenuation. In contrast, $DDPG$ and $TD3$ maintain relatively stable SRS values, indicating their effectiveness in leveraging semantic information. Notably, SAMRAMARL outperforms both traditional algorithms, showcasing its advanced resource allocation strategies.
Figure 5: semantic demand size v.s. QOE
Fig. 5 illustrates the impact of semantic demand size on the QoE. As the size of the semantic data packets increases, the QoE for all algorithms improves, reflecting the direct relationship between semantic demand size and QoE. Larger semantic data packets allow more comprehensive information transmission, enhancing overall communication effectiveness. However, at a semantic demand size of 2000 bytes per $u_{m}$, SAMRAMARL achieves a QoE level that is relatively mid-range among the three methods. At a demand size of 3000 bytes per $u_{m}$, we observe a decline in QoE for $DDPG$ and $TD3$, indicating their struggle to adapt to the increased data load. In contrast, SAMRAMARL retains stable performance, which can be attributed to its robust reinforcement learning framework.
Figure 6: Semantic demand size v.s. SRS
Fig. 6 examines the relationship between semantic demand size and SRS. The curve for $DDPG_{N}O_{S}C$ declines linearly as demand size increases, highlighting its inability to manage the growing volume of semantic data. In contrast, $DDPG$ and $TD3$ show stability, effectively utilizing semantic context to maintain performance. SAMRAMARL consistently achieves higher SRS, attributed to its multi-agent reinforcement learning framework, which facilitates resource distribution and adapts to vehicular environment nuances.
Figure 7: Transform factor ${u_{m}}$ v.s. QOE
Fig. 7 shows the relationship between the transform factor ${u_{m}}$ and QoE. The $DDPG_{N}O_{S}C$ curve declines with increasing ${u_{m}}$, reflecting the traditional QoE calculation’s sensitivity to complexity. However, semantic-aware algorithms maintain stable QoE levels, as they focus on the relevance of transmitted information rather than just symbol length. Among these, SAMRAMARL achieves the highest QoE due to its dynamic resource allocation capabilities.
Figure 8: Transform factor ${u_{m}}$ v.s. SRS
@@ V Conclusion
  • While QoE initially increases with semantic demand sizes, a threshold is identified beyond which QoE stabilizes, emphasizing the need for effective semantic information handling to avoid degradation in performance.

### v2

@@ III-A Modeling of the System
State space: The state of the $n$th PL at time slot $t$ consists of the following components: the instantaneous channel information of the $n$th PL and the BS $h_{n,BS}^{t}[k]$, the instantaneous channel information of the $n$th PL and his followers $h_{n,m}^{t}[k]$, the previous interference from other platoons $I_{n,m}^{t}[k]$ and $I_{T,BS}^{t}[k]$, the residuals transmitted by the intra-platoon communication loads $B_{s,n}^{t}$, and thus, the state space can be expressed as:
$s_{n}^{t}=\left[h_{n,BS}^{t}[k],h_{n,m}^{t}[k],I_{n,m}^{t}[k],I_{T,BS}^{t}[k],B_{s,n}^{t}\right].$ (24)
Action space: As mentioned before, PL contains several actions: $\beta_{n,k}$ is used to select the sub-channel for communication, $\rho_{n,k}$ determines inter or intra communication, the power level of the communication $p_{T,n}[k]$ and $p_{I,n}[k]$, and the sentence lengths corresponding to V2I communication and V2V communication ($u_{m}^{T}$ and $u_{m}^{I}$). Note that when the PL selects inter communication, it has only one sentence length to select; when the PL selects intra communication, there will be $M$ followers of sentence lengths to select. Thus, the action space of agent $n$ at time slot $t$ is defined as:
@@
Reward function: in our proposed scheme, there are two aspects of rewards for each agent, one is the global reward reflecting the cooperation among agents, and the other is the local reward that is to help each agent to explore the optimal action. In DRL, rewards can be set flexibly, and a good reward can improve the performance of the system. Here, our goal is to optimize the agent $n$’s QoE level, as well as the corresponding payload transmission probability. Thus the local and global rewards of the $n$th PL at time slot $t$ are:
$\begin{array}[]{l}{r_{n,l}^{t}}=-{w_{1}}\frac{1}{{1+\exp(-\alpha(\sum\limits_{t=1}^{T}{\sum\limits_{m=1}^{M_{n}}{{\beta_{n,k}}}}\phi_{k}[m,t]-\frac{{{B_{s}}}}{{\Delta T}}))}}\\ {\rm{}}\qquad\qquad\qquad\qquad\qquad\qquad\quad+{w_{2}}Qo{E^{n}},\end{array}$ (26)
@@ III-B SAMRAMARL Algorithm
Algorithm 1 SAMRAMARL Algorithm
The SAMRAMARL algorithm we proposed is designed to optimize both global system objectives and individual agent behavior in a dynamic, semantic communication environment, particularly tailored for vehicular platooning scenarios. The algorithm employs a multi-agent actor-critic approach, where each agent is equipped with its own actor and local critic networks for decision-making, while a pair of global critics guides the overall system optimization [24].
Specifically, we consider a vehicular environment with $N$ platoons and the policies for all agents are $\mathbf{\pi}=\{{\pi_{1}},{\pi_{2}},\cdots,{\pi_{N}}\}$. The agent $n$’s strategy $\pi_{n}$, Q-functions ${\mathop{\rm Q}\nolimits}_{{\phi_{n}}}^{k}$ and twin global critic Q-functions (${\mathop{\rm Q}\nolimits}_{\psi 1}^{{g_{1}}}$, ${\mathop{\rm Q}\nolimits}_{\psi 2}^{{g_{2}}}$) are parameterized by ${\theta_{n}}$, $\phi_{n}$, ${\psi_{1}}$ and $\psi_{2}$, respectively. The training process begins with an initial the environment simulator, configured with $P$ vehicular platoons to reflect a realistic multi-agent communication scenario. Initially, twin global critic networks, $Q_{{\psi_{1}}}^{{g_{1}}}$ and $Q_{{\psi_{2}}}^{{g_{2}}}$ are initialized to evaluate joint state-action pairs across all agents. These global critics are responsible for assessing the performance of the entire system based on global rewards which is QOE. Alongside, each agent $n$ initializes its policy network ${\pi_{n}}$ for action selection and a local critic network $Q_{{\phi_{n}}}^{n}$ to evaluate the quality of its actions based on local observations.
@@
The detailed SAMRAMARL algorithm is described in the Algorithm 1. Since the algorithm is implemented as the actor-critic framework, denote the computational complexities for calculating gradients and updating parameters in the actor and critic networks being denoted as $G_{A}$, $G_{C}$, $U_{A}$, and $U_{C}$ respectively. Since the architecture of the target actor network and the target critic network is the same as the network structure of the two mentioned above, their complexities remain the same. Thus, the computational complexity of the SAMRAMARL algorithm is calculated as $\mathcal{O}\left({{G_{A}}+{G_{C}}+2{U_{A}}+2{U_{C}}}\right)+{\text{ }}\mathcal{O}\left({2{G_{C}}+4{U_{C}}}\right)$.
@@ IV-B Performance Evaluation
To assess the performance of the proposed distributed semantic-aware multi-modal resource allocation framework (SAMRAMARL), we utilize QoE and SRS in V2X links as performance metrics. Furthermore, communication delay is also considered, since it can directly reflect the real-time performance of semantic information transmission. We compare SAMRAMARL against the following baseline methods:
@@
In traditional communication systems, semantic information is transmitted using bits rather than symbols explicitly designed for semantic representation. Although each bit can be loosely considered a semantic symbol, it generally conveys less semantic information compared to symbols employed in semantic transmission methods like DeepSC. Therefore, in order to compare with semantic communication, $QOE^{\prime}$ is used in traditional communication, defined similar to equation (15):
@@
where ${u_{m}^{T/I}}$ is defined as the transform factor, with unit of bits/word. For example, if a word consists of 5 letters and ASCII code is used for encoding, then ${u_{m}^{T/I}}$ is 40 bits/word. We assume that ${{\tilde{H}}_{SM/MM}}$ is fixed, indicating no bit errors in conventional communications.
@@
In Fig. 2, we analyze the reward evolution of various algorithms, including our SAMRAMARL framework. SAMRAMARL exhibits faster convergence and greater stability than $DDPG$ and $TD3$, attributable to its multi-agent structure. This decentralized decision-making allows each vehicle to respond independently to local conditions, optimizing resource allocation in dynamic environments. While $DDPG$ and $TD3$ also consider semantic awareness, their single-agent models struggle to adapt quickly to environmental changes. Specifically, $DDPG$ faces instability from a poor exploration-exploitation balance, resulting in fluctuating rewards, while $TD3$ converges more slowly than SAMRAMARL, highlighting the benefits of multi-agent systems. The distributed nature of SAMRAMARL fosters better collaboration, leading to quicker convergence and stability, particularly under rapidly changing semantic information. This approach effectively meets the varying demands of vehicles within a platoon, ensuring efficient spectrum sharing and resource allocation.
Figure 3: Intra platoon Gap v.s. QOE with semantic demand size $=4$ ksut.
@@
Figure 4: Intra platoon Gap v.s. SRS with semantic demand size $=4$ ksuts.
Fig. 4 presents the relationship between intra-platoon gap and SRS. The $DDPG\_NO\_SC$ curve shows a clear decline with increasing gaps, reflecting increased signal attenuation. In contrast, $DDPG$ and $TD3$ maintain relatively stable SRS values, indicating their effectiveness in leveraging semantic information. Notably, SAMRAMARL outperforms both traditional algorithms, showcasing its advanced resource allocation strategies.
Figure 5: semantic demand size v.s. QOE with intra platoon gap $=20$ m.
Fig. 5 illustrates the impact of semantic demand size on the QoE. As the size of the semantic data packets increases, the QoE for all algorithms improves, reflecting the direct relationship between semantic demand size and QoE. Larger semantic data packets allow more comprehensive information transmission, enhancing overall communication effectiveness. However, at a semantic demand size of 2000 suts, SAMRAMARL achieves a QoE level that is relatively mid-range among the three methods. At a demand size of 3000 suts, we observe a decline in QoE for $DDPG$ and $TD3$, indicating their struggle to adapt to the increased data load. In contrast, SAMRAMARL retains stable performance, which can be attributed to its robust reinforcement learning framework.
Figure 6: Semantic demand size v.s. SRS with intra platoon gap $=20$ m.
Fig. 6 examines the relationship between semantic demand size and SRS. The curve for $DDPG\_NO\_SC$ declines linearly as demand size increases, highlighting its inability to manage the growing volume of semantic data. In contrast, $DDPG$ and $TD3$ show stability, effectively utilizing semantic context to maintain performance. SAMRAMARL consistently achieves higher SRS, attributed to its multi-agent reinforcement learning framework, which facilitates resource distribution and adapts to vehicular environment nuances.
Figure 7: Transform factor ${u_{m}}$ v.s. QOE with semantic demand size $=4$ ksut and intra platoon gap $=20$ m.
Fig. 7 shows the relationship between the transform factor ${u_{m}}$ and QoE. The $DDPG\_NO\_SC$ curve declines with increasing ${u_{m}}$, reflecting the traditional QoE calculation’s sensitivity to complexity. However, semantic-aware algorithms maintain stable QoE levels, as they focus on the relevance of transmitted information rather than just symbol length. Among these, SAMRAMARL achieves the highest QoE due to its dynamic resource allocation capabilities.
Figure 8: Transform factor ${u_{m}}$ v.s. SRS with semantic demand size $=4$ ksut and intra platoon gap $=20$ m.
@@
Figure 9: Semantic demand size ${u_{m}}$ v.s. Delay with intra platoon gap $=20$ m.
In Fig. 9, the communication delay increases for all methods as the semantic demand size increases, but communication delay of SAMRAMARL is always the lowest and has the smallest growth rate. This is attributed to the multi-agent reinforcement learning framework, which can efficiently manage the transmission of semantic information and significantly enhance the robustness of the system. $DDPG\_NO\_SC$ has a relatively large delay because it cannot perceive semantics. Moreover, with the increase of semantic demand size, the delay of $DDPG\_NO\_SC$ shows an obvious growth trend, indicating that it cannot effectively cope with the load pressure brought by the increase of semantic demand size. $DDPG$ and $TD3$ are superior to $DDPG\_NO\_SC$, because they can sense semantics and significantly reduce transmission redundancy, thereby reducing delay. The performance of $TD3$ in terms of delay is similar to that of $DDPG$, but the overall delay is slightly lower due to the double-delay mechanism of $TD3$.
@@ V Conclusion
  • While QoE initially increases with semantic demand sizes, a threshold is identified beyond which QoE stabilizes, emphasizing the need for effective semantic feature information handling to avoid degradation in performance.